You borrowed $10000000 in the commercial paper market. The rate on 30-day commercial paper is 6.48 percent.
(a) If you issued 30-day commercial paper through a dealer, how much must you repay at the end of the 30-day period?
(b) Assume that, 30 days from now, CP rates have not changed. Instead of repaying your borrowings, you want to roll them over for another 30 days. How much mustyou repay for the end of the second 30 days?
Answer #1

Assumptions: Interest Rate is Annual rate

a)

Repayment at the end of 30 days = Amount Borrowed*[1+(Interest Rate*30/365)] = 10000000*[1+(0.0648*30/365)] = 10000000*1.005326

= $10053260.274

b)

Repayment at the end of NEXT 30 days = Repayment at the end of FIRST 30 days*[1+(Interest Rate*30/365)] = 10053260.274*[1+(0.0648*30/365)] = 10053260.274*1.005326

= $10106803.938